# Rectified Flows for Fast Multiscale Fluid Flow Modeling

### v1

@@ Appendix A Architecture of the ReFlow Model
Figure A.1: UNet Backbone Architecture Used in ReFlow. This schematic illustrates the core UNet-based architecture used within the ReFlow framework, structured across three resolution levels. For clarity, the number of blocks per level is set to one in this illustration. In the actual GenCFD and ReFlow configurations, each block is repeated four times per level. The bottleneck shows an asymmetry between the encoder and decoder sides: the block on the encoder side includes a Convolution Block, Positional Embedding, and Attention Block, while the corresponding block on the decoder side omits the Positional Embedding.
A.1 Architectural Components
@@
Architecture Layout. We employ a UNet-based architecture augmented with attentional layers, typically termed as UViT by [28]. In addition to the MLP-based time embedding $\Gamma(\tau)$, the input signal is lifted into a higher-dimensional embedding space and later projected back into physical space through convolutional layers. The data flows from high to low resolution across a down-sampling stack composed of convolution blocks into a bottleneck layer that encodes the fine-scale features of the inputs across its channel dimension and mixes them via a multi-head attention mechanism.
@@
Importantly, the architecture described here resembles the one presented in [21], ensuring consistency with prior state-of-the-art practices in high-fidelity generative modeling for scientific computing.
A.1.1 Convolution Block
@@
A.1.2 Normalized FiLM Layer
@@
$\mathbf{z}$ $=\;\mathrm{Conv}\bigl{(}\mathrm{Norm}(\mathbf{x})\bigr{)},\qquad\mathbf{z}\;=\;\sigma(\mathbf{z}+\alpha\,\gamma(t)),$ (A.1)
@@ A.1.3 Attention Block
$\mathbf{Q}\;=\;W_{Q}\,\mathbf{h},\quad\mathbf{K}\;=\;W_{K}\,\mathbf{h},\quad\mathbf{V}\;=\;W_{V}\,\mathbf{h},\quad\mathcal{A}(\mathbf{h})\;=\;\mathrm{softmax}\!\Bigl{(}\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\Bigr{)}\,\mathbf{V},$ (A.2)
@@
$\mathrm{MHA}(\mathbf{h})\;=\;W^{O}\;\mathrm{Concat}\left(\mathcal{A}_{1}(\mathbf{h}),\dots,\mathcal{A}_{n}(\mathbf{h})\right),$ (A.3)
@@
$\mathbf{z}$ $=\;\mathrm{MHA}\bigl{(}\mathrm{Norm}(\mathbf{x})\bigr{)},$ (A.4)
@@
A.2 Training Pipeline Details
@@
Table A.1: Representative hyperparameters for 2D fluid-flow tasks. See main text for justification of these choices.
@@
ODE Steps (Sampling) 4–8 Often >15×>15$\times$> 15 × fewer vs. diffusion.
@@
In practice we noticed that the most robust denoising schedule was just the uniform one. For details on the cos-map schedule, we direct the reader to its description in [3].
A.3 Summary and Key Takeaways
The RectifiedFlow class, combined with a time-aware UNet2D architecture, builds a deterministic transport model well-suited to multi-scale PDE tasks. By interpolating between noise and data along nearly-straight paths, it reduces sampling overhead compared to conventional diffusion PDE solvers while retaining the flexibility to model complex turbulent phenomena. Further experimental and theoretical results appear in SM Section D. Instead of the multi-step stochastic integration in diffusion-based models, RectifiedFlow fits a deterministic ODE from noise to data. This leverages straighter trajectories in frequency space, enabling fewer network evaluations at inference time.
@@ A.3 Summary and Key Takeaways
We solve this with odeint. Few steps ($8\!-\!10$) typically suffice, contrasting with the tens or hundreds required by full diffusion samplers.
Appendix B Datasets
B.1 Multi-scale Flow Datasets
Following the dataset generation procedure of Herde et al. [11], the Richtmyer–Meshkov ensemble is created by imposing a randomized sinusoidal perturbation on a two-fluid interface and driving it with a planar shock via prescribed pressure jumps. For the CloudShock and ShearFlow benchmarks, we follow the GenCFD mesh-generation and solver setup of Molinaro et al. [21], initializing CloudShock with concentric density perturbations and ShearFlow with orthogonal shear jets on adaptive Cartesian grids, and employing the same high-order finite-volume scheme and dissipation settings to ensure consistent numerical fidelity across both datasets.
Overview. We experiment on three paired-field benchmarks $(u_{0},u_{1})$:
@@ B.1 Multi-scale Flow Datasets
  • Shear-Layer (SL): 79 200 samples
@@
Macro–Micro Ensemble Evaluation (CS & SL). To probe out-of-distribution generalization under small perturbations, we adopt a two-stage ensemble protocol:
@@
    Macro-sampling: select $M_{\mathrm{macro}}=10$ distinct base initial conditions.
@@
    Micro-perturbation: for each base, generate $M_{\mathrm{micro}}=1{,}000$ perturbed copies within a small radius.
@@
    Metrics: compute all performance measures (mean-field error $e_{\mu}$, standard-deviation error $e_{\sigma}$, average $W_{1}$, spectral agreement) by first averaging over each micro-ensemble, then averaging across the $M_{\mathrm{macro}}$ cases.
In-Distribution Evaluation (RM). For the Richtmyer–Meshkov dataset, we additionally evaluate on a held-out test set drawn from the same distribution (no macro–micro perturbations). This allows us to study how deterministic and probabilistic models scale with training set size under standard, in-distribution conditions.
B.2 Preprocessing and Data Organization
@@
Before training, we compute and store global, channel-wise statistics over the entire training dataset. Concretely, for each physical variable $c$ we calculate
@@
where $N_{\rm train}$ is the number of training samples, $H\times W\,[\times D]$ the spatial grid size, and $u_{c}(x,y,[z],t_{0})$ the initial field for channel $c$. We average over all spatial locations to yield a single mean $\mu_{c}$ and standard deviation $\sigma_{c}$ per channel. The same statistics are also computed over the final-time fields $u_{c}(t_{f})$, producing $\mu_{c}^{\prime}$ and $\sigma_{c}^{\prime}$ for the output channels.
@@
where $u_{c}$ and $v_{c}$ denote the initial and target fields, respectively. We then concatenate $\{\tilde{u}_{c}\}$ and $\{\tilde{v}_{c}\}$ along the channel axis before feeding them to the network. Because we always apply the same pre-computed statistics at test time, no information from the evaluation set ever influences these normalization parameters.
Test-time ensemble perturbation (evaluation only).
We adopt the same testing setup proposed in GenCFD [21]. To evaluate how well our model captures the distribution of solutions $u(t)$ arising from a fixed, chaotic initial field $\bar{u}$, we generate a small perturbed ensemble around $\bar{u}$ and evolve each member with a high-fidelity PDE solver—but only at evaluation time. Specifically:
@@
    Micro-ensemble generation. Around the test initial condition $\bar{u}$, draw $M_{\mathrm{micro}}$ perturbed copies $\{\bar{u}_{j}\}_{j=1}^{M_{\mathrm{micro}}}$ by sampling uniformly in a ball of radius $\varepsilon$ centered at $\bar{u}$.
@@
    Reference propagation. Integrate each perturbed field $\bar{u}_{j}$ forward to time $t$ using the reference solver, producing end-states $\{u_{j}(t)\}$.
@@
When the input distribution itself is non-degenerate, one would first draw $M_{\mathrm{macro}}$ base fields from that distribution and then apply the micro-ensemble procedure to each. For our Dirac initial-condition tests we set $M_{\mathrm{macro}}=1$.
@@
During training the model only ever sees independent pairs $(u_{0},u_{1})$—one target per input—and never observes any ensemble. The above ensemble perturbation is used only to construct a ground-truth distribution at test time.
Appendix C Models and baselines
@@
Shared Backbone. We stress that for GenCFD, GenCFD $\circ$ FNO and UViT we use the exact same UViT backbone architecture outlined in the earlier parameterization of our rectified flow as in Table A.1 for the sake of consistency. The architecture is further illustrated schematically in Figure A.1, which shows the UNet-based design in detail.
@@
GenCFD [21] is an end-to-end conditional score-based diffusion model that directly learns the mapping from an input flow field $u_{0}$ to the target field $u_{1}$. It uses the same UViT backbone described in Table A.1, with noise levels $\sigma_{\tau}$ corresponding to the standard deviation of the added Gaussian perturbation, in contrast to our ReFlow model which evolves over a rectified diffusion time $\tau\in[0,1]$. Common choices for noise schedulers include exponential and tangent formulations. In our case, we consistently adopted an exponential noise schedule, as it yielded the best results across all datasets and metrics reported. For score-based diffusion models, either variance-exploding (VE) or variance-preserving (VP) formulations are typically used. In all GenCFD models, we opted for the VE setting throughout. During training we minimize the denoising score-matching loss over $(u_{0},u_{1})$ pairs sampled uniformly in $\tau\in[0,1]$, using Adam with initial learning rate $3\times 10^{-4}$ (cosine decay), batch size 16, and EMA decay 0.9999. As is common in score-based diffusion models, we applied denoiser preconditioning and used uniform weighting across noise levels. At inference we solve the corresponding SDE (via an Euler-Maruyama scheme with discretization $128$ steps) and incorporate reconstruction guidance to condition on $u_{0}$ so that we recover the conditional posterior measure $p(u_{1}|u_{0})$.
GenCFD $\boldsymbol{\circ}$ FNO
In this hybrid approach, a Fourier Neural Operator (FNO) [14] is first trained under an $\ell_{2}$ loss to predict $u_{1}$ from $u_{0}$. Its output $\hat{u}_{1}^{\rm FNO}$ is concatenated channel-wise with $u_{0}$ and fed into the GenCFD score network at both train and test time. All other architectural and optimization settings (backbone, scheduler, SDE sampler) remain identical to GenCFD, allowing the FNO to provide fast low-frequency priors while the diffusion stage refines fine-scale turbulent features.
Custom implementation.This schematic closely parallels the pipeline of [23], but with two key distinctions. First, we fully replace both the diffusion backbone and its denoising stages with GenCFD. Second, and most important, we condition not only on the FNO’s coarse solution but also on the original high-resolution initial condition. This ensures that our model retains fine-scale structures that would otherwise be lost if we relied solely on the FNO predictions.
@@
Our FNO baseline follows Li et al. [14] and is implemented in PyTorch. First, each spatial location of the input field $u_{0}$ is lifted from $C_{\rm in}$ to 256 channels via a two-layer MLP with a SiLU activation, then projected down to a 128-channel hidden representation. This is followed by four FnoResBlocks: in each block, the hidden features undergo a Conditional LayerNorm (time embedding size 128), SiLU, and two spectral convolutions (SpectralConv) with $num\_modes=(24,24)$, combined through a learnable “soft-gate” residual skip. Finally, a two-layer projection MLP (128 → 256 → $C_{\rm out}$, with SiLU in between) produces the output field $\hat{u}_{1}$. We train for 100k iterations minimizing $\|u_{1}-\hat{u}_{1}\|_{2}^{2}$ with Adam (learning rate $10^{-3}$, batch size 16), and perform inference in one forward pass—no diffusion or autoregression is used.
@@
Our UViT surrogate [28] uses the identical UNet2D backbone of Table A.1, injecting time embeddings and employing 8-head multi-head attention at the bottleneck. It is trained deterministically with MSE loss on $(u_{0},u_{1})$, using Adam at $3\times 10^{-4}$, batch size 16, for 100 k iterations (EMA 0.9999). For multi-step forecasting we roll out autoregressively, feeding each prediction back as the next input. This ablates out the diffusion component while matching all other design choices.
@@
Table C.1 summarizes key architectural settings for all baseline models considered in this work. These configurations were selected to reflect the strongest performance achievable within a shared training budget of up to 120,000 iterations, as used in the quantitative comparison reported in Table  5.1.
@@
Table C.1: Summary of architectural depth and approximated parameter count for each model evaluated. All UNet-based models (ReFlow, GenCFD, GenCFD $\circ$ FNO, and UViT) share a consistent backbone design to ensure consistency across diffusion and non-diffusion baselines, while the FNO baseline uses a structurally different architecture. These configurations reflect the best-performing settings under a training budget of up to 120,000 iterations, as used in the statistical comparisons reported in Table 5.1.
@@
Appendix D Results
D.1 Performance Metrics
@@
  • Mean-field error. Let $\mu_{\rm ref}(x)$ and $\mu_{\rm model}(x)$ denote the pointwise spatial means of the reference and model ensembles, respectively. We measure their discrepancy by the $L^{2}$-norm
@@
  • Standard-deviation error. Similarly, let $\sigma_{\rm ref}(x)$ and $\sigma_{\rm model}(x)$ be the pointwise standard deviations. We therefore compute
@@
  • Average 1-Wasserstein distance. For a fixed initial condition $u_{0}$, let $\{u^{(j)}_{\rm ref}(x)\}_{j=1}^{N}$ and $\{u^{(j)}_{\rm model}(x)\}_{j=1}^{N}$ be the ensembles of final-time values produced by the reference solver and our model, respectively. Since these fields are discretized on $M$ spatial points $\{x_{i}\}_{i=1}^{M}$, we obtain two empirical 1D distributions at each $x_{i}$. Denoting their inverse CDFs by $F^{-1}_{\rm ref}(\cdot\,;\,x_{i})$ and $F^{-1}_{\rm model}(\cdot\,;\,x_{i})$, the pointwise Wasserstein-1 distance is
@@
Energy spectra. Finally, to assess how well different scales are captured, we compute the discrete energy spectrum of each velocity field. Given a sample $u(x)$ on a $d$-dimensional periodic grid with spacing $\Delta$, let $\widehat{u}_{k}$ be its Fourier coefficient at integer wavevector $k\in\mathbb{Z}^{d}$. We bin modes by their $\ell_{1}$ radius $|k|_{1}=k_{1}+\dots+k_{d}$, defining
@@
and compare the ensemble-average spectra of reference and model. Our plots show $E_{r}$ versus $r$ on log–log axes to highlight scale-by-scale agreement.
D.2 Training Efficiency Observations
@@
Figure D.1 summarizes performance over the training trajectory. At approximately 100,000 steps, ReFlow consistently outperforms GenCFD across all key metrics: relative $L^{2}$ error in the mean ($e_{\mu}$), relative $L^{2}$ error in the standard deviation ($e_{\sigma}$), and the average pointwise Wasserstein-1 (W1) distance—evaluated for both $u_{x}$ and $u_{y}$ velocity components. Notably, the $y$-axis uses a logarithmic scale to emphasize differences in error magnitude across training steps.
@@
To further probe the long-term capabilities of GenCFD, we extended its training to 600,000 iterations and checked whether it eventually surpasses ReFlow’s performance at its final recorded checkpoint (150,000 steps). Table D.1 summarizes the earliest training steps at which GenCFD overtakes ReFlow for each evaluation metric. While GenCFD does catch up and exceed ReFlow in several cases, it never surpasses ReFlow on the Wasserstein-1 distance for $u_{y}$ within the training horizon considered.
Table D.1: Number of training iterations at which GenCFD surpasses ReFlow’s performance at 150,000 iterations. If not surpassed, indicated by “—”.

### v2

@@ Appendix I From law-level bounds to reported pointwise metrics
Figure 5: UNet Backbone Architecture Used in ReFlow. This schematic illustrates the core UNet-based architecture used within the ReFlow framework, structured across three resolution levels. For clarity, the number of blocks per level is set to one in this illustration. In the actual GenCFD and ReFlow configurations, each block is repeated four times per level. The bottleneck shows an asymmetry between the encoder and decoder sides: the block on the encoder side includes a Convolution Block, Positional Embedding, and Attention Block, while the corresponding block on the decoder side omits the Positional Embedding.
J.1 Architectural Components
@@
Architecture Layout. We employ a UNet-based architecture augmented with attentional layers, typically termed as UViT by Saharia et al. (2022). In addition to the MLP-based time embedding $\Gamma(\tau)$, the input signal is lifted into a higher-dimensional embedding space and later projected back into physical space through convolutional layers. The data flows from high to low resolution across a down-sampling stack composed of convolution blocks into a bottleneck layer that encodes the fine-scale features of the inputs across its channel dimension and mixes them via a multi-head attention mechanism.
@@
Importantly, the architecture described here resembles the one presented in Molinaro et al. (2024), ensuring consistency with prior state-of-the-art practices in high-fidelity generative modeling for scientific computing.
J.1.1 Convolution Block
@@
J.1.2 Normalized FiLM Layer
@@
$\mathbf{z}$ $=\;\mathrm{Conv}\bigl(\mathrm{Norm}(\mathbf{x})\bigr),\qquad\mathbf{z}\;=\;\sigma(\mathbf{z}+\alpha\,\gamma(t)),$ (35)
@@
$\mathbf{Q}\;=\;W_{Q}\,\mathbf{h},\quad\mathbf{K}\;=\;W_{K}\,\mathbf{h},\quad\mathbf{V}\;=\;W_{V}\,\mathbf{h},\quad\mathcal{A}(\mathbf{h})\;=\;\mathrm{softmax}\!\Bigl(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\Bigr)\,\mathbf{V},$ (36)
@@
$\mathrm{MHA}(\mathbf{h})\;=\;W^{O}\;\mathrm{Concat}\left(\mathcal{A}_{1}(\mathbf{h}),\dots,\mathcal{A}_{n}(\mathbf{h})\right),$ (37)
@@
$\mathbf{z}$ $=\;\mathrm{MHA}\bigl(\mathrm{Norm}(\mathbf{x})\bigr),$ (38)
@@
J.2 Training Pipeline Details
@@
Table 5: Representative hyperparameters for 2D fluid-flow tasks. See main text for justification of these choices.
@@
ODE Steps (Sampling) 4–8 Often $15\times$ fewer vs. diffusion.
@@
In practice we noticed that the most robust denoising schedule was just the uniform one. For details on the cos-map schedule, we direct the reader to its description in Esser et al. (2024).
J.3 Summary and Key Takeaways
The RectifiedFlow class, combined with a time-aware UNet2D architecture, builds a deterministic transport model well-suited to multi-scale PDE tasks. By interpolating between noise and data along nearly-straight paths, it reduces sampling overhead compared to conventional diffusion PDE solvers while retaining the flexibility to model complex turbulent phenomena. Further experimental and theoretical results appear in SM Section M. Instead of the multi-step stochastic integration in diffusion-based models, RectifiedFlow fits a deterministic ODE from noise to data. This leverages straighter trajectories in frequency space, enabling fewer network evaluations at inference time.
@@
We solve this with odeint. Few steps (8108\!-\!10) typically suffice, contrasting with the tens or hundreds required by full diffusion samplers.
Appendix K Datasets
K.1 Multi-scale Flow Datasets
Following the dataset generation procedure of Herde et al. Herde et al. (2024), the Richtmyer–Meshkov ensemble is created by imposing a randomized sinusoidal perturbation on a two‐fluid interface and driving it with a planar shock via prescribed pressure jumps. For the CloudShock and ShearFlow benchmarks, we follow the GenCFD mesh‐generation and solver setup of Molinaro et al. Molinaro et al. (2024), initializing CloudShock with concentric density perturbations and ShearFlow with orthogonal shear jets on adaptive Cartesian grids, and employing the same high‐order finite‐volume scheme and dissipation settings to ensure consistent numerical fidelity across both datasets.
Overview. We experiment on three paired‐field benchmarks $(u_{0},u_{1})$:
@@
  • Shear‐Layer (SL): 79 200 samples
@@
Macro–Micro Ensemble Evaluation (CS & SL). To probe out‐of‐distribution generalization under small perturbations, we adopt a two‐stage ensemble protocol:
@@
    Macro‐sampling: select $M_{\mathrm{macro}}=10$ distinct base initial conditions.
@@
    Micro‐perturbation: for each base, generate $M_{\mathrm{micro}}=1{,}000$ perturbed copies within a small radius.
@@
    Metrics: compute all performance measures (mean‐field error $e_{\mu}$, standard‐deviation error $e_{\sigma}$, average $W_{1}$, spectral agreement) by first averaging over each micro‐ensemble, then averaging across the $M_{\mathrm{macro}}$ cases.
In‐Distribution Evaluation (RM). For the Richtmyer–Meshkov dataset, we additionally evaluate on a held‐out test set drawn from the same distribution (no macro–micro perturbations). This allows us to study how deterministic and probabilistic models scale with training set size under standard, in‐distribution conditions.
K.2 Preprocessing and Data Organization
@@
Before training, we compute and store global, channel‐wise statistics over the entire training dataset. Concretely, for each physical variable $c$ we calculate
@@
where $N_{\rm train}$ is the number of training samples, $H\times W\,[\times D]$ the spatial grid size, and $u_{c}(x,y,[z],t_{0})$ the initial field for channel $c$. We average over all spatial locations to yield a single mean $\mu_{c}$ and standard deviation $\sigma_{c}$ per channel. The same statistics are also computed over the final‐time fields $u_{c}(t_{f})$, producing $\mu_{c}^{\prime}$ and $\sigma_{c}^{\prime}$ for the output channels.
@@
where $u_{c}$ and $v_{c}$ denote the initial and target fields, respectively. We then concatenate $\{\tilde{u}_{c}\}$ and $\{\tilde{v}_{c}\}$ along the channel axis before feeding them to the network. Because we always apply the same pre‐computed statistics at test time, no information from the evaluation set ever influences these normalization parameters.
Test‐time ensemble perturbation (evaluation only).
We adopt the same testing setup proposed in GenCFD Molinaro et al. (2024). To evaluate how well our model captures the distribution of solutions $u(t)$ arising from a fixed, chaotic initial field $\bar{u}$, we generate a small perturbed ensemble around $\bar{u}$ and evolve each member with a high‐fidelity PDE solver—but only at evaluation time. Specifically:
@@
    Micro‐ensemble generation. Around the test initial condition $\bar{u}$, draw $M_{\mathrm{micro}}$ perturbed copies $\{\bar{u}_{j}\}_{j=1}^{M_{\mathrm{micro}}}$ by sampling uniformly in a ball of radius $\varepsilon$ centered at $\bar{u}$.
@@
    Reference propagation. Integrate each perturbed field $\bar{u}_{j}$ forward to time $t$ using the reference solver, producing end‐states $\{u_{j}(t)\}$.
@@
When the input distribution itself is non‐degenerate, one would first draw $M_{\mathrm{macro}}$ base fields from that distribution and then apply the micro‐ensemble procedure to each. For our Dirac initial‐condition tests we set $M_{\mathrm{macro}}=1$.
@@
During training the model only ever sees independent pairs $(u_{0},u_{1})$—one target per input—and never observes any ensemble. The above ensemble perturbation is used only to construct a ground‐truth distribution at test time.
Appendix L Models and baselines
@@
Shared Backbone. We stress that for GenCFD, GenCFD $\circ$ FNO and UViT we use the exact same UViT backbone architecture outlined in the earlier parameterization of our rectified flow as in Table 5 for the sake of consistency. The architecture is further illustrated schematically in Figure 5, which shows the UNet-based design in detail.
@@
GenCFD Molinaro et al. (2024) is an end-to-end conditional score‐based diffusion model that directly learns the mapping from an input flow field $u_{0}$ to the target field $u_{1}$. It uses the same UViT backbone described in Table 5, with noise levels $\sigma_{\tau}$ corresponding to the standard deviation of the added Gaussian perturbation, in contrast to our ReFlow model which evolves over a rectified diffusion time $\tau\in[0,1]$. Common choices for noise schedulers include exponential and tangent formulations. In our case, we consistently adopted an exponential noise schedule, as it yielded the best results across all datasets and metrics reported. For score-based diffusion models, either variance-exploding (VE) or variance-preserving (VP) formulations are typically used. In all GenCFD models, we opted for the VE setting throughout. During training we minimize the denoising score‐matching loss over $(u_{0},u_{1})$ pairs sampled uniformly in $\tau\in[0,1]$, using Adam with initial learning rate $3\times 10^{-4}$ (cosine decay), batch size 16, and EMA decay 0.9999. As is common in score-based diffusion models, we applied denoiser preconditioning and used uniform weighting across noise levels. At inference we solve the corresponding SDE (via an Euler-Maruyama scheme with discretization $128$ steps) and incorporate reconstruction guidance to condition on $u_{0}$ so that we recover the conditional posterior measure $p(u_{1}|u_{0})$.
[…]
In this hybrid approach, a Fourier Neural Operator (FNO) Li et al. (2020a) is first trained under an $\ell_{2}$ loss to predict $u_{1}$ from $u_{0}$. Its output $\hat{u}_{1}^{\rm FNO}$ is concatenated channel‐wise with $u_{0}$ and fed into the GenCFD score network at both train and test time. All other architectural and optimization settings (backbone, scheduler, SDE sampler) remain identical to GenCFD, allowing the FNO to provide fast low‐frequency priors while the diffusion stage refines fine‐scale turbulent features.
Custom implementation.This schematic closely parallels the pipeline of Oommen et al. (2024), but with two key distinctions. First, we fully replace both the diffusion backbone and its denoising stages with GenCFD. Second, and most important, we condition not only on the FNO’s coarse solution but also on the original high-resolution initial condition. This ensures that our model retains fine-scale structures that would otherwise be lost if we relied solely on the FNO predictions.
@@
Our FNO baseline follows Li et al. Li et al. (2020a) and is implemented in PyTorch. First, each spatial location of the input field $u_{0}$ is lifted from $C_{\rm in}$ to 256 channels via a two‐layer MLP with a SiLU activation, then projected down to a 128‐channel hidden representation. This is followed by four FnoResBlocks: in each block, the hidden features undergo a Conditional LayerNorm (time embedding size 128), SiLU, and two spectral convolutions (SpectralConv) with num_modes=(24,24)num\_modes=(24,24), combined through a learnable “soft‐gate” residual skip. Finally, a two‐layer projection MLP (128 → 256 → $C_{\rm out}$, with SiLU in between) produces the output field $\hat{u}_{1}$. We train for 100k iterations minimizing $\|u_{1}-\hat{u}_{1}\|_{2}^{2}$ with Adam (learning rate $10^{-3}$, batch size 16), and perform inference in one forward pass—no diffusion or autoregression is used.
@@
Our UViT surrogate Saharia et al. (2022) uses the identical UNet2D backbone of Table 5, injecting time embeddings and employing 8‐head multi‐head attention at the bottleneck. It is trained deterministically with MSE loss on $(u_{0},u_{1})$, using Adam at $3\times 10^{-4}$, batch size 16, for 100 k iterations (EMA 0.9999). For multi‐step forecasting we roll out autoregressively, feeding each prediction back as the next input. This ablates out the diffusion component while matching all other design choices.
@@
Table 6 summarizes key architectural settings for all baseline models considered in this work. These configurations were selected to reflect the strongest performance achievable within a shared training budget of up to 120,000 iterations, as used in the quantitative comparison reported in Table  1.
@@
Table 6: Summary of architectural depth and approximated parameter count for each model evaluated. All UNet-based models (ReFlow, GenCFD, GenCFD $\circ$ FNO, and UViT) share a consistent backbone design to ensure consistency across diffusion and non-diffusion baselines, while the FNO baseline uses a structurally different architecture. These configurations reflect the best-performing settings under a training budget of up to 120,000 iterations, as used in the statistical comparisons reported in Table 1.
@@
Appendix M Results
M.1 Performance Metrics
@@
  • Mean‐field error. Let $\mu_{\rm ref}(x)$ and $\mu_{\rm model}(x)$ denote the pointwise spatial means of the reference and model ensembles, respectively. We measure their discrepancy by the $L^{2}$‐norm
@@
  • Standard‐deviation error. Similarly, let $\sigma_{\rm ref}(x)$ and $\sigma_{\rm model}(x)$ be the pointwise standard deviations. We therefore compute
@@
  • Average 1-Wasserstein distance. For a fixed initial condition $u_{0}$, let $\{u^{(j)}_{\rm ref}(x)\}_{j=1}^{N}$ and $\{u^{(j)}_{\rm model}(x)\}_{j=1}^{N}$ be the ensembles of final‐time values produced by the reference solver and our model, respectively. Since these fields are discretized on $M$ spatial points $\{x_{i}\}_{i=1}^{M}$, we obtain two empirical 1D distributions at each $x_{i}$. Denoting their inverse CDFs by $F^{-1}_{\rm ref}(\cdot\,;\,x_{i})$ and $F^{-1}_{\rm model}(\cdot\,;\,x_{i})$, the pointwise Wasserstein‐1 distance is
@@
Energy spectra. Finally, to assess how well different scales are captured, we compute the discrete energy spectrum of each velocity field. Given a sample $u(x)$ on a dd‐dimensional periodic grid with spacing $\Delta$, let $\widehat{u}_{k}$ be its Fourier coefficient at integer wavevector $k\in\mathbb{Z}^{d}$. We bin modes by their $\ell_{1}$ radius $|k|_{1}=k_{1}+\dots+k_{d}$, defining
@@
and compare the ensemble‐average spectra of reference and model. Our plots show $E_{r}$ versus $r$ on log–log axes to highlight scale‐by‐scale agreement.
M.2 Training Efficiency Observations
@@
Figure 6 summarizes performance over the training trajectory. At approximately 100,000 steps, ReFlow consistently outperforms GenCFD across all key metrics: relative $L^{2}$ error in the mean ($e_{\mu}$), relative $L^{2}$ error in the standard deviation ($e_{\sigma}$), and the average pointwise Wasserstein-1 (W1) distance—evaluated for both $u_{x}$ and $u_{y}$ velocity components. Notably, the $y$-axis uses a logarithmic scale to emphasize differences in error magnitude across training steps.
@@
To further probe the long-term capabilities of GenCFD, we extended its training to 600,000 iterations and checked whether it eventually surpasses ReFlow’s performance at its final recorded checkpoint (150,000 steps). Table 7 summarizes the earliest training steps at which GenCFD overtakes ReFlow for each evaluation metric. While GenCFD does catch up and exceed ReFlow in several cases, it never surpasses ReFlow on the Wasserstein-1 distance for $u_{y}$ within the training horizon considered.
Table 7: Number of training iterations at which GenCFD surpasses ReFlow’s performance at 150,000 iterations. If not surpassed, indicated by “—”.